Need help with these shelf-life calculations!! Step by step would be great! Thanks

1. The specific decomposition rate k of an experimental drug obtained from Arrhenius plot is   3.4 x 10-4 hr-1 at room temperature. If the time required for 10% of a drug to degrade (t90%) represents a reasonable limit of degradation for the active ingredient and is designated as shelf-life, what will be the shelf-life in hours of this drug:

2. The shelf life of an API in a solution was found to be 25 days. What is the drug
Answer #1

1. The specific decomposition rate k of an experimental drug obtained from Arrhenius plot is 3.4 x 10-4 hr-1 at room temperature. If the time required for 10% of a drug to degrade (t90%) represents a reasonable limit of degradation for the active ingredient and is designated as shelf-life, what will be the shelf-life in hours of this drug:

Arrhenius plot

k = 3.4 x 10-4 hr-1

Reaction is of first order from the unit of k.

10% degradation, 90%A remains.

A/A0 = 0.9

Ln(A0/A) = k*t90

Shelf-life time t90 = Ln(1/0.9)/(3.4*10^-4) = 310hr

2. The shelf life of an API in a solution was found to be 25 days. What is the drug